Suppose production function is
Q=k²L²-0.12k²-3.2L²
Then find the level of output where
APL=MPL,if k=15
MPK=0,if L=10
1
Expert's answer
2020-08-17T10:08:53-0400

1.

"APL=\frac{Q}{L}"


"MPL=\frac{\Delta Q}{\Delta L}"


"Q=k^2L^2-0.12k^2-3.2L^2=15^2L^2-0.12\times15^2-3.2L^2=225L^2-27-3.2L^2=221.8L^2-27"


put Q into the formula:


"L=\sqrt{\frac{Q+0.12k^2}{k^2-3.2}}=\sqrt{\frac{Q+27}{221.8}}=\sqrt{L^2-\frac{27}{221.8}}"


APL=MPL


"Q=APL\times L"


put the found expressions in the formula and get Q


L cut on each other


Q=221.8L^2-27


"\Delta Q=APL\times\Delta L"

"\Delta Q=\frac{Q}{L}\times\Delta L"


"\Delta Q=\frac{221.8L^2-27}{\sqrt{L^2-\frac{27}{221.8}}}\times(\frac{1}{2\times \sqrt{L^2-\frac{27}{221.8}}})=110.9"

2.


"MPK=\frac{\Delta Q}{\Delta k}"


MPK=0 L=10


"\Delta Q=MPK\times\Delta k"


"\Delta Q=0"


Find the derivatives Q


Q'=2k2L-0.24k-6.4L=0

"2k2\times10-0.24k-6.4\times10=0"

40k-0.24k-64=0

39.76k=64

"k=\frac{64}{39.76}"

"Q'=2\frac{64}{39.76}2\times10-0.24\frac{64}{39.76}-6.4\times10=0"